OpenStudy (anonymous):

What is the discontinuity and zero of the function f(x) = 2x^2+5x-12/x+4? Discontinuity at (-4, -11), zero at (3/2 , 0) Discontinuity at (-4, -11), zero at (- 3/2 , 0) Discontinuity at (4, 5), zero at ( 3/2, 0) Discontinuity at (4, 5), zero at (-3/2 , 0)

OpenStudy (anonymous):

Anybody know? @Kainui

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@MathLegend @mathmale

OpenStudy (anonymous):

@Mertsj

OpenStudy (anonymous):

@wio

OpenStudy (mertsj):

Can the denominator be 0?

OpenStudy (anonymous):

no those are the responses up above! @Mertsj

OpenStudy (anonymous):

i just really need help with this problem! @Mertsj @wio

OpenStudy (anonymous):

So it would be out of the list of answer choices it would be either the first or second one! @Mertsj

OpenStudy (mertsj):

first one

OpenStudy (anonymous):

thank you!

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

@Mertsj Here is another one if you could help me with it!

OpenStudy (anonymous):

What are the zero(s) of the function f(x)=4x^2-36x/x-9

OpenStudy (mertsj):

Factor the numerator

OpenStudy (anonymous):

@Mertsj answer choices! x = -9 x = 0 x = 9 x = 0 and x = 9

OpenStudy (mertsj):

\[\frac{4x^2-36x}{x-9}\]

OpenStudy (mertsj):

Is that the problem?

OpenStudy (anonymous):

yup!

OpenStudy (mertsj):

Factor the numerator

OpenStudy (anonymous):

Elaborate @Mertsj

OpenStudy (mertsj):

The numerator is 4x^2-36x Factor out 4x

OpenStudy (anonymous):

x^2 -9

OpenStudy (mertsj):

4x(x^2-9) Does that = 4x^2-36x?

OpenStudy (anonymous):

yes? @Mertsj

OpenStudy (mertsj):

What is 4x multiplied by x^2

OpenStudy (anonymous):

4x^3 @Mertsj

OpenStudy (mertsj):

Then I guess those could not be the factors because there is on 4x^3 in your problem so factor the numerator and this time makes sure the factors multiply back

OpenStudy (anonymous):

Um i dont understand? can you explain it to me? @Mertsj

OpenStudy (mertsj):

You are working on a pre calc test. You need to discover on your own how to factor out the common factor from 4x^2-36x.

OpenStudy (mertsj):

That is an algebra 1 problem

OpenStudy (anonymous):

isint it 0 and 9

OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):

The denominator is x-9 so x cannot be 9

OpenStudy (anonymous):

Then it is 0 @Mertsj

OpenStudy (anonymous):

?? @Mertsj

OpenStudy (anonymous):

correct @Mertsj

OpenStudy (tkhunny):

Please note that the original problem: f(x) = 2x^2+5x-12/x+4? You have written: \(f(x) = 2x^{2} + 5x - \dfrac{12}{x} + 4\) If that is not your intent, please remember your Order of Operations and use parentheses as necessary.

OpenStudy (tkhunny):

Distributive Property of Multiplication over Addition. a(b+c) = ab + ac or the other way ab + ac = a(b+c) \( 4x^2-36x = 4(x^{2} - 9x)\) Do the same with the constant factor "x".

OpenStudy (anonymous):

ok but wouldnt that cause for the answers to be equal to 0 or no?

OpenStudy (tkhunny):

* common factor, not constant factor. Sorry about that.

OpenStudy (tkhunny):

I don't actually care what the answers are, just yet. I want to SEE it factored and WRITTEN. Then we can worry about the final solution.

OpenStudy (anonymous):

Oh ok I am sorry! @tkhunny